OpenStudy (anonymous):

Find two numbers to make the equation true. 3m2 = 147

OpenStudy (zzr0ck3r):

\[2m^2=147\rightarrow m^2 = 147/3 = 49 \rightarrow m=\sqrt{49} \rightarrow m=\pm? \]

OpenStudy (anonymous):

So um whats the answer

OpenStudy (zzr0ck3r):

what is the square root of 49?

OpenStudy (anonymous):

oh ok thanks